# Radio over Fiber with Cascaded Structure: Algorithm for Uplink Positioning

### v1

@@ IV-A RU with PAs operating in the linear regime
Under this scenario, the whole RS system is linear time-invariant. Thus, the signal model considers the dispersive channel of PMFs and the wireless channel from the UE to the entry RU. It is assumed that the UE transmits a known sequence of samples, $\boldsymbol{s}=[s_{0},\cdots,s_{K-1}]^{\mathrm{T}}\in\mathbb{C}^{K\times 1}$ over $K$ subcarriers. Initially, the signals undergo the wireless channel between the UE and the RU, which is unknown to the CU. Due to the sparse nature of the sub-THz wireless channel, the line-of-sight (LoS) components are considered to be dominant. Thus, the $k$th element of the input signal at the RoF, $S_{k}$ with $\boldsymbol{x}=[x_{0},x_{1},\cdots,x_{K-1}]^{\mathrm{T}}$ (see Fig. 1), can be written as
@@
where
  • AA is a complex value absorbing the effects of the wireless channel and the phase offset between the UE and the RoF. The magnitude of $A$ represents the pathloss in LoS.
  • $\tau$ is the unknown propagation delay including TOA and the clock offset, which can be expressed as
    $\tau=\frac{d}{c}+\delta_{\text{t}}$ (7)
    with $c$ being the speed of light, and $\delta_{\text{t}}$ being the clock offset between the UE and the RoF.
Note that distinguishing the effects of $d$ and $\delta_{\text{t}}$ is not possible since they have the same influence over the signals, i.e. phase shift and delay (see (7)). However, one can rely on the effects of cascaded fibers and PAs to estimate $r$ and $\tau$ separately.
The $k$th element of the output signal at the entry RU denoted as $y_{k}^{(1)}$ is given by
@@
with $w^{(r)}_{k}\sim\mathcal{CN}(0,\sigma^{2})$ being the $k$th AWGN noise component at the $r$th RU. At the output of the next RU, after propagating through one segment of dispersive fiber and one PA, the signal at $k$th frequency is given by
$y_{k}^{(2)}=G^{2}H_{k}Ae^{-j2\pi f_{k}\tau}s_{k}+GH_{k}w_{k}^{(0)}+w_{k}^{(1)}.$ (9)
In a general form, after propagating through $r$ RUs along the RoF, the received signal, with $\boldsymbol{y}=[y_{0},y_{1},\cdots,y_{K-1}]^{\mathrm{T}}$, at the CU can be written as
$y_{k}^{(r)}=G^{(r+1)}H_{k}^{r}Ae^{-j2\pi f_{k}\tau}s_{k}+\underbrace{G^{(r)}H_{k}^{r}w_{k}^{(0)}+\cdots+w_{k}^{(r)}}_{(r+1)\,\text{independent noises}}.$ (10)
Note that the cumulative filtered noise by linear PAs and fibers still follow the Gaussian distribution since the linear transformation does not change the probability distribution. Let $\tilde{w}_{k}$ denote the sum of independent Gaussian noise terms, then the variance of $\tilde{w}_{k}$ is given by
@@
where $b_{k}\triangleq(G\lvert H_{k}\rvert)^{2}.$ Until now, parameter $r$ takes discrete values, indicating the number of stages that the sub-THz signals propagate in a RoF. Since CRLB only applied to continuous-valued variables, we propose a continuous-value representation of $r$ in Proposition 1.
@@
Note that (14) only holds for $b_{k}\neq 1,\forall k$. In other words, there are two cases worth being studied. Here, we name them as flat fiber and frequency selective fiber cases.
@@ IV-A1 Flat fiber
Let us consider a special case in which the frequency response of the fiber is totally flat, that is the case for a system with very limited bandwidth. With the use of PMF, the frequency response can be very flat at 1 GHz bandwidth. It is assumed that the amplification $G$ is set to compensate for the attenuation, corresponding to $b_{k}=1,\forall k$. Consequently, the variance of effective noise is
@@
and the phase shift induced by the fiber grows linearly with the propagation distance $r$. In other words, a flat fiber only supports one-mode transmission within the fiber, corresponding to a single-tap impulse response.
@@
IV-A2 Frequency selective fiber
A frequency selective fiber means the magnitude of its frequency response fluctuates over the bandwidth of interest, which makes $b_{k}$ not equal to one. Without the loss of generality, we assume $b_{k}\neq 1,\forall k$, which ensures the validity of (14) for all frequencies. This is the case for a system with a large bandwidth, for example, 10 GHz. Consequently, the variance of effective noise is
@@ IV-A2 Frequency selective fiber
and the phase shifts introduced by the fiber might grow non-linearly with the propagation distance $r$. Similarly, one can view a frequency selective fiber to support multi-mode transmission, whose corresponding impulse response has multiple taps.
@@ IV-B RU with PAs operating in the non-linear regime
For EE purposes, the PAs are allowed to work in the non-linear regime, as modeled in Section III. As previously stated, the UE transmits a known sequence of samples $\boldsymbol{s}=[s_{0},\cdots,s_{K-1}]^{\mathrm{T}}$ over $K$ subcarriers to the RoF. We model the input signal in the time domain as
@@
To obtain the output at the following RUs, we define the function $f(\cdot)$, that combines the effects of a non-linear PA and one segment of PMF, and can be expressed as
@@
Therefore, the received signal at the CU can be expressed as $\boldsymbol{y}^{(r)}=f^{r}\left(\boldsymbol{y}^{(0)}\right)+\boldsymbol{w}^{(r)}$, where all the unknown parameters are interwoven with each other. As a result, we cannot apply the same estimator proposed in the linear regime ((19) and (20)). Alternatively, we can employ a NLS framework.
With PAs operating in the non-linear regime, a NLS framework may not serve as an optimal estimator as the noise distribution is transformed from a Gaussian distribution into another probability distribution. However, the output of a NLS framework will not deviate significantly from the optimum. Following the NLS framework, the objective function w.r.t. the unknown parameter $A,r,$ and $\tau$ is given by
@@ V Cramér-Rao Lower Bound
For simplicity, we express the complex valued $A$ and $H$ with their magnitudes $\lvert A\rvert$ and $\lvert H_{k}\rvert$ and phases $\phi$ and $\psi_{k}$, i.e., $A=\lvert A\rvert e^{j\phi}$ and $H=\lvert H_{k}\rvert e^{j\psi_{k}}$. Hence, the unknown parameters are defined as $\boldsymbol{\theta}\triangleq[\lvert A\rvert,\phi,\tau,r]^{\mathrm{T}}.$
The definition for the CRLB is given by
@@
which reveals the amount of information associated with each pair of unknown parameters. One important characteristic of FIM is that every FIM is positive semi-definite, indicating the information content is non-negative and ensuring its eigenvalues are non-negative. Moreover, a positive definite FIM indicates the parameters are identifiable and a singular FIM indicates the parameters are non-identifiable.
As mentioned in Section IV, there are two expressions for the variance of effective noise as shown in (15) and in (21). Naturally, the CRLB needs to be calculated for both flat fiber and frequency selective fiber cases. Detailed calculations can be found in the Appendix Radio over Fiber with Cascaded Structure: Algorithm for Uplink Positioning.
@@ V-A CRLB for flat fiber
Therefore, partial derivatives in (29) can be calculated as
$\left[\frac{\partial\boldsymbol{\mu(\theta)}}{\partial\lvert A\rvert}\right]_{k}=$ $Ge^{j(\phi+r\psi_{k}-2\pi f_{k}\tau)}s_{k},$ (31a)
$\left[\frac{\partial\boldsymbol{\mu(\theta)}}{\partial\phi}\right]_{k}=$ $jG\lvert A\rvert s_{k}e^{j(\phi+r\psi_{k}-2\pi f_{k}\tau)},$ (31b)
$\left[\frac{\partial\boldsymbol{\mu(\theta)}}{\partial\tau}\right]_{k}=$ $-j2\pi f_{k}G\lvert A\rvert s_{k}e^{j(\phi+r\psi_{k}-2\pi f_{k}\tau)},$ (31c)
$\left[\frac{\partial\boldsymbol{\mu(\theta)}}{\partial r}\right]_{k}=$ $G\lvert A\rvert e^{j(\phi+r\psi_{k}-2\pi f_{k}\tau)}s_{k}\left(j\psi_{k}\right).$ (31d)
According to the symmetry property, $\left[\boldsymbol{I}(\theta)\right]_{i,j}=\left[\boldsymbol{I}(\theta)\right]_{j,i}$ is always valid so that there are 10 unique entries in each FIM. Regarding the flat fiber, the elements of the FIM are calculated as
@@
(a) Group delay
(b) Magnitude
Figure 6: Measured characteristics of a 1-meter PMF made of high-density polyethylene with a solid rectangular cross-section in the D-band (110 GHz to 170 GHz) [13]: (a) Group delay; (b) Magnitude.
@@
    According to (32i), the system can rely on the overall noise variance to estimate the propagation distance, as it accumulates linearly with $r$.
V-B CRLB for frequency selective fiber
For the frequency selective fiber case, the noise variance is defined as
@@ V-B CRLB for frequency selective fiber
and $\boldsymbol{\mu(\theta)}$ is the same as (30). One can derive the first-order derivatives as
$\times\left(\!\ln\!{\sqrt{b_{k}}}\!+\!j\psi_{k}\!\right)\!.$ (34d)
Similarly, the entries in the FIM can be obtained as
@@
Discussion: Based on the CRLB obtained for the frequency selective fiber, one can observe
@@
    With the use of frequency selective fiber, a coupling between $\lvert A\rvert$ and $r$ appears as illustrated in (35d), which comes from the varied magnitude of the frequency response of the PMF.
@@ VI Simulation results
In this section, we first validate the proposed NLS estimator, relying on the measured channel of a segment of PMF. Then, we validate the proposed CRLB with two artificial channels generated under a constraint in the total energy.
@@ VI-A Validation of the proposed estimators with measured fiber channel characteristics
In this section, we illustrate the performance obtained by the proposed estimators with the fiber channel measurements presented in [13]. These measurements inherently include several imperfections, such as noise, interference, and measurement errors. Therefore, a median filtering with a window size of 300 is implemented through Matlab to smooth the data in group delay and magnitude, separately. Both the smoothed data and the original data are shown in Figs. 6(a) and 6(b). The measured data consists of channel characteristics at $K$ discrete frequencies, encompassing both magnitudes ($\lvert\boldsymbol{H}\rvert=\big[\lvert H_{0}\rvert,\cdots,\lvert H_{K-1}\rvert\big]^{\mathrm{T}}$) and group delay. Subsequently, the phase response $\boldsymbol{\psi}=[\psi_{0},\cdots,\psi_{K-1}]^{\mathrm{T}}$ was obtained by integrating the group delay on the discrete frequencies. Then, the transfer function can be expressed as
$\boldsymbol{H}=\lvert\boldsymbol{H}\rvert\circ\exp{(j\boldsymbol{\psi)}},$ (36)
where $\circ$ represents the Hadamard product. The employed LOS pathloss model comprises free-space path loss and a large-scale fading component, which can be written as
@@
with $\lambda$ being the wavelength, $d$ being defined in (1), $G_{t}$ and $G_{r}$ being the transmitter gain, and receiver gain respectively. The large-scale fading component $\zeta$ follows a log-normal distribution with a standard deviation of 2 dB. This model can be replaced by other developed LoS models for sub-THz propagation,such as the 3GPP model in [24].
@@
Monte Carlo simulations were performed with a RoF system comprising five RUs and one CU. Between every two components, a fiber with the frequency response as expressed in (36) serves as the connection. Distortion incurred by PAs was illustrated in Fig. 4. The UE transmits a known sequence of samples using a quadrature phase shift keying waveform with a unit amplitude. The UE accesses the RoF system through the third RU. The channel magnitude illustrated in Fig. 6(b) corresponds to a spectrum with 1 GHz bandwidth centered around 140 GHz, from which the maximum magnitude in a one-meter fiber is approximately -2.48 dB. Therefore, $G$ is set to 2.48 dB for PAs. The parameters in (37) were set to ensure the input amplitude is in a certain regime (linear or non-linear) of the PAs.
The proposed PSO was implemented in Matlab and the results are shown in Fig. 7. It is observed that the error rate would decrease to a level of $10^{-5}$ with a proper input amplitude, which indicates that the proposed algorithm is able to accurately estimate the propagation distance even when PAs work in the non-linear regime. Through comparison between the estimation accuracy with two non-linear factors, one can observe an accuracy loss with a higher non-linear factor.

### v2

@@ IV-A RU with PAs operating in the linear regime
Following the discussion in Section III, this paper considers PAs working in both linear and non-linear regimes. In this subsection, let us consider PAs working in the linear regime. Under this scenario, the whole RoF system is LTI. Thus, the signal model considers the dispersive channel of PMFs and the wireless channel from the UE to the entry RU. It is assumed that the UE transmits a known sequence of samples, $\boldsymbol{s}=[s_{0},\cdots,s_{K-1}]^{\mathrm{T}}\in\mathbb{C}^{K\times 1}$ over $K$ frequency points. Initially, the signals undergo the wireless channel between the UE and the RU, which is unknown to the CU.
Thus, the $k$th element of the input signal at the RoF, $x_{k}$ with $\boldsymbol{x}=[x_{0},x_{1},\cdots,x_{K-1}]^{\mathrm{T}}$ (see Fig. 1), can be written as
@@
where $\tau$ can be expressed as
$\tau=\frac{d}{c}+\delta_{\text{t}},$ (9)
with $c$ being the speed of light, and $\delta_{\text{t}}$ being the clock offset between the UE and the RoF. Note that distinguishing the effects of $d$ and $\delta_{\text{t}}$ is not possible since they have the same influence over the signals, i.e., phase shift and delay (see (9)). However, one can rely on the effects of cascaded fibers and PAs to estimate $r$ and $\tau$ separately.
The $k$th element of the output signal at the entry RU, denoted as $y_{k}^{(0)}$, is given by
@@
with $w^{(r)}_{k}\sim\mathcal{CN}(0,\sigma^{2})$ being the $k$th AWGN noise component at the $r$th stage. The wireless channel as such does not contribute noise; only the PAs add noise, as they are the only active electronic components.
In a general form, after propagating through $r$ units (i.e., $r$ PMF segments propagation distance) along the RoF, the received signal at the CU, with $\boldsymbol{y}^{(r)}=[y_{0}^{(r)},y_{1}^{(r)},\cdots,y_{K-1}^{(r)}]^{\mathrm{T}}$, can be written as
$y_{k}^{(r)}=G^{r+1}H_{k}^{r}Ae^{-j2\pi f_{k}\tau}s_{k}+\underbrace{G^{r}H_{k}^{r}w_{k}^{(0)}+\cdots+w_{k}^{(r)}}_{(r+1)\,\text{independent noises}}.$ (11)
Note that the CU also amplifies the signals. Hence, there are $(r+1)$ PAs and $r$ PMF segments.
The cumulative filtered noise by linear PAs and fibers still follows the Gaussian distribution since the linear transformation does not change the probability distribution. Let $\tilde{w}_{k}$ denote the sum of independent Gaussian noise terms, then the variance of $\tilde{w}_{k}$ is given by
@@
where $b_{k}\triangleq(G\lvert H_{k}\rvert)^{2}.$ Until now, the parameter $r$ takes discrete values, indicating the number of stages that the sub-THz signals propagate in an RoF stripe. Since CRLB only applied to continuous-valued variables, we propose a continuous-value representation of $r$ in Proposition 1.
@@
Note that (15) only holds for $b_{k}\neq 1,\forall k$. In other words, there are two cases worth being studied. We name them as flat fiber and frequency-selective fiber cases. This approach was also adopted in [Henk_GLOBCOM_2023] to calculate the CRLB of a discrete variable.
@@ IV-A1 Flat fiber
Let us consider a special case in which the frequency response of the fiber is totally flat, that is, the case for a system with very limited bandwidth. With the use of PMF, the frequency response can be very flat at 1 GHz bandwidth. It is assumed that the amplification $G$ is set to compensate for the attenuation, corresponding to $b_{k}=1,\forall k$. Consequently, the variance of effective noise is
@@
and the phase shift induced by the fiber grows linearly with the propagation distance $r$. In other words, a flat fiber only supports one-mode transmission within the fiber, corresponding to a single-tap impulse response. It should be noted that the system can not separate the effects of the wireless channel and the RoF based on the phase, but only relies on the noise variance shown in (16) to estimate $r$.
@@
IV-A2 Frequency-selective fiber
A frequency-selective fiber means the magnitude of its frequency response fluctuates over the bandwidth of interest, which makes $b_{k}$ not equal to one. Without the loss of generality, we assume $b_{k}\neq 1,\forall k$, which ensures the validity of (15) for all frequencies. This is the case for a system with a large bandwidth, for example, 10 GHz. Consequently, the variance of effective noise is
@@ IV-A2 Frequency-selective fiber
and the phase shifts introduced by the fiber might grow non-linearly with the propagation distance $r$. Similarly, one can view a frequency-selective fiber to support multi-mode transmission, whose corresponding impulse response has multiple taps.
@@ IV-B RU with PAs operating in the non-linear regime
For EE purposes, the PAs are allowed to work in the non-linear regime, as modeled in Section III. As previously stated, the UE transmits a known sequence of samples $\boldsymbol{s}=[s_{0},\cdots,s_{K-1}]^{\mathrm{T}}$ over $K$ frequency points to the RoF. We model the input signal in the time domain as
@@
To obtain the output at the following units, we define the function $f(\cdot)$, which combines the effects of a non-linear PA and one segment of PMF, and can be expressed as
@@
Therefore, the received signal at the CU can be expressed as $\boldsymbol{y}^{(r)}=f^{r}\left(\boldsymbol{y}^{(0)}\right)+\boldsymbol{w}^{(r)}$, where all the unknown parameters are interwoven with each other. As a result, we cannot apply the same estimator proposed in the linear regime ((20) and (21)). Alternatively, we can employ an NLS framework.
With PAs operating in the non-linear regime, an NLS framework may not serve as an optimal estimator as the noise distribution is transformed from a Gaussian distribution into another probability distribution. However, the output of an NLS framework will not deviate significantly from the optimum. Following the NLS framework, the objective function w.r.t. the unknown parameter $A,r,$ and $\tau$ is given by
@@
IV-C Positioning of the UE
According to (5), the distance between the $i$th UE and the entry RUs is written as $\boldsymbol{d}\triangleq[d_{1,i},d_{2,i},d_{3,i}]^{\mathrm{T}}$. The ToA from the UE to the entry RUs is interwoven with a clock offset as defined in (9), which can be viewed as a bias. The position of the UE $U_{x}^{(i)},U_{y}^{(i)}$ is estimated through the estimated $\boldsymbol{\tau}=(1/c)\boldsymbol{d}+\mathbf{1}_{3}\delta_{t}$ at three RoF stripes and can be written as
$\begin{bmatrix}\hat{U_{x}}^{(i)}\\ \hat{U_{y}}^{(i)}\end{bmatrix}=\mathop{\min}_{U_{x}^{(i)},U_{y}^{(i)},\delta_{\text{t}}}\lvert\lvert\boldsymbol{\tau}-\frac{\boldsymbol{d}}{c}-\mathbf{1}_{3}\delta_{\text{t}}\rvert\rvert^{2},$ (29)
$\begin{bmatrix}\hat{U_{x}}^{(i)}\\ \hat{U_{y}}^{(i)}\end{bmatrix}=\mathop{\min}_{U_{x}^{(i)},U_{y}^{(i)}}\lvert\lvert\boldsymbol{R}-\mathbf{1}_{3}(\mathbf{1}_{3}^{\mathrm{T}}\mathbf{1}_{3})^{-1}\mathbf{1}_{3}^{\mathrm{T}}\boldsymbol{R}\rvert\rvert^{2}.$ (31)
IV-D The effect of the non-ideal PA
The analysis of this section is based on ideal PA assumptions, i.e., all PAs have identical and constant characteristics. Although this assumption is ideal, a small deviation from it does not lead to significant effects on the estimators. Practically, the noise variance and PA gain show little variation with the room temperature (much less than 2 dB gain variation in [zhou2022chip]). The cumulative impact of cascading PAs is much more dominant than that of varying individual PA characteristics. For example, two PAs in a cascade doubles the noise variance and squares the gain, compared to a single PA.
@@ V Cramér-Rao Lower Bound
For simplicity, we express the complex valued $A$ and $H$ with their magnitudes $\lvert A\rvert$ and $\lvert H_{k}\rvert$ and phases $\phi$ and $\psi_{k}$, i.e., $A=\lvert A\rvert e^{j\phi}$ and $H=\lvert H_{k}\rvert e^{j\psi_{k}}$. Hence, the unknown parameters are defined as $\boldsymbol{\theta}\triangleq[\lvert A\rvert,\phi,\tau,r]^{\mathrm{T}}.$ The definition for the CRLB is given by
@@
which reveals the amount of information associated with each pair of unknown parameters. The FIM in (33) is valid only for Gaussian models. One important characteristic of FIM is that every FIM is positive semi-definite, indicating the information content is non-negative and ensuring its eigenvalues are non-negative. Moreover, a positive definite FIM indicates the parameters are identifiable, and a singular FIM indicates the parameters are non-identifiable.
As mentioned in Section IV, there are two expressions for the variance of effective noise as shown in (16) and in (22). Naturally, the CRLB needs to be calculated for both flat fiber and frequency-selective fiber cases. Detailed calculations can be found in the Appendix Acknowledgment.
@@ V-A CRLB for flat fiber
According to the symmetry property, $\left[\boldsymbol{I}(\theta)\right]_{i,j}=\left[\boldsymbol{I}(\theta)\right]_{j,i}$ is always valid so that there are 10 unique entries in each FIM. Regarding the flat fiber, the elements of the FIM are
@@
Figure 7: Measured characteristics of a 1-meter PMF made of high-density polyethylene with a solid rectangular cross-section in the D-band (110 GHz to 170 GHz) [Frida]: (a) Group delay; (b) Magnitude.
@@
    According to (35i), the system can rely on the overall noise variance to estimate the propagation distance, as it cumulates linearly with $r$.
V-B CRLB for frequency-selective fiber
For the frequency-selective fiber case, the noise variance is defined as
@@ V-B CRLB for frequency-selective fiber
and $\boldsymbol{\mu(\theta)}$ is the same as (34). Similarly, the entries in the FIM can be obtained as
@@
Discussion: Based on the CRLB obtained for the frequency-selective fiber, one can observe
@@
    With the use of frequency-selective fiber, a coupling between $\lvert A\rvert$ and $r$ appears as illustrated in (37d), which comes from the varied magnitude of the frequency response of the PMF.
@@ VI Simulation results
In this section, we first validate the proposed NLS estimator, relying on the measured channel of a segment of PMF. Then, we validate the proposed CRLB with two channels generated under a constraint on the total energy. Moreover, we apply the proposed positioning algorithm in an indoor scenario.
@@ VI-A Validation of the proposed estimators with measured fiber channel characteristics
In this section, we illustrate the performance obtained by the proposed estimators with the fiber channel measurements presented in [Frida]. The channel measurement data in [Frida] was obtained by transmitting sub-THz (D-band) signals over a one-meter long high-density polyethylene PMF with a solid rectangular cross-section of 11 $\times 2 mm$ and a density of $0.93$ g/cm3. These measurements inherently include several imperfections, such as noise, interference, and measurement errors. Therefore, a median filtering with a window size of 300 is implemented through MATLAB to smooth the data in group delay and magnitude, separately. Both the smoothed data and the original data are shown in Figs. 7(a) and 7(b).
The measured data consists of channel characteristics at $K$ discrete frequencies, encompassing both magnitudes ($\lvert\boldsymbol{H}\rvert=\big[\lvert H_{0}\rvert,\cdots,\lvert H_{K-1}\rvert\big]^{\mathrm{T}}$) and group delay. Subsequently, the phase response $\boldsymbol{\psi}=[\psi_{0},\cdots,\psi_{K-1}]^{\mathrm{T}}$ was obtained by integrating the group delay on the discrete frequencies. Then, the transfer function can be expressed as
$H_{k}=\lvert H_{k}\rvert\exp{(j\psi_{k})}.$ (38)
The employed LoS model comprises free-space path loss and a large-scale fading component, which can be written as
@@
with $\lambda$ being the wavelength, $d$ being defined in (5), $G_{t}$ and $G_{r}$ being the transmitter gain, and receiver gain respectively. The large-scale fading component $\zeta$ follows a log-normal distribution with a standard deviation of 2 dB. This model can be replaced by other developed LoS models for sub-THz propagation, such as the 3GPP model in [3gpp]. In our simulation, $G_{t}$ is 36 dB, and $G_{r}$ is also 36 dB. Assuming the use of state-of-the-art patch antennas with 9 dBi gain [Yuyan_Eucap_2024, Yuyan_Eucap_2025], a 22 $\times$ 22 array is required, which translates to 4 cm $\times$ 4 cm antenna size. A horn antenna with equivalent gain can also be used.
@@
Monte Carlo simulations were performed with an RoF system comprising five RUs and BUs, and one CU. Between every two units, a fiber with the frequency response as expressed in (38) serves as the connection. Distortion incurred by PAs was illustrated in Fig. 4. The UE transmits a known sequence of samples using a quadrature phase shift keying waveform with a unit amplitude. The UE accesses the RoF system through the third RU. The channel response shown in Fig. 7(b) corresponds to a spectrum with 1 GHz bandwidth centered around 140 GHz, from which the magnitude in a one-meter fiber is approximately -2.48 dB. Therefore, $G$ is set to 2.48 dB for PAs. The parameters in (39) were set to ensure the input amplitude is in a certain regime (linear or non-linear) of PAs.
The proposed NLS estimator was implemented in Matlab, and the results are shown in Fig. 8. It is observed that the error rate would decrease to a level of $10^{-5}$ with a proper input amplitude, which indicates that the proposed algorithm is able to accurately estimate the propagation distance even when PAs work in the non-linear regime. Through comparison between the estimation accuracy with two non-linear factors, one can observe an accuracy loss with a higher non-linear factor.